# Holographic Correlators of Giant Gravitons

### v1

@@ 1 Introduction
As we will show, the problem of studying holographic correlators for giant gravitons boils down to studying charged geodesics hanging from the boundary in the 5d gauged SUGRA. This is essentially the WKB approximation to the two-point function [17]. It turns out that, in the presence of the monodromy defect, there are two relevant such geodesics. One is the direct relative of the U-shaped geodesic capturing the two-point function in the absence of the defect. The other is a novel saddle corresponding to a geodesic anchored to the defect whose existence is tied to the presence of the defect. The full two-point function is the sum of both contributions, with the anchored contribution capturing the one-point function for the operator $OO^{\dagger}$ in the presence of the defect. It is important to emphasize that while the particular coefficients will depend on details of the defect, the structure and presence of the contribution itself is universal, only controlled by the existence of the geodesic and the asymptotics of the background.
@@ 2 Giant gravitons as charged geodesics in 5d gauged supergravity
Thus we see that effectively our D3 brane behaves like a charged particle, with charge equals to the mass, in the 5d geometry, precisely as expected for a giant graviton. The mass $m$ coincides with the dimension of the giant graviton and charge equaling the mass reflects the BPS condition saturated by the giant. We emphasize that (8) shows that $m\sim N$.
@@ 5 Comments on the anchored geodesics
From here it follows that, provided $\beta_{1}\neq 0$, the anchored geodesic looks like the cartoon in figure 1\@vpageref[]fig:anchored.
@@
On the other hand, this non-analytic behavior in a two-point function is strange from the standpoint of finite $N$ QFT. Thus, it is natural to expect the existence of a mechanism to smooth this $\Theta$ behavior. One possible such mechanism is suggested by inspection of figure 1\@vpageref[]fig:anchored, since, as $\beta_{1}\rightarrow 0$ the configuration looks like a brane/antibrane prone to decay through nucleation of D3-brane tubes breaking the anchored geodesic. A cartoon of the configuration would be as in figure 2\@vpageref[]fig:instability below.
@@
$S_{\rm D3-tube}=m\frac{P^{\frac{1}{2}}}{H^{\frac{2}{3}}}\,\Delta\theta\,.$ (49)
@@
$S_{\rm D3-tube}=m\frac{(P^{\prime}_{\star})^{\frac{1}{2}}}{H_{\star}^{\frac{2}{3}}}\,\sqrt{y-y_{\star}}\,\Delta\theta\sim m\,\sqrt{y-y_{\star}}\,\Delta\theta\,.$ (50)
Close to $y_{\star}$, since we want the tube to meet the two arms of the U, we can use the equation of motion and write $\sqrt{y-y_{\star}}\,\Delta\theta=\frac{\beta_{1}}{2}$, [^6] This explains in retrospective why we did not include a Chern-Simons interaction: close to $y_{\star}$ the CS part would have been roughly $\beta_{1}\,(y-y_{\star})\,\Delta\theta\sim\beta_{1}^{2}$. so we can write
$S_{\rm D3-tube}\sim\frac{m\beta_{1}}{2}\,.$ (51)
Hence we find a constant action for a tube located in a region close but not exactly equal to $y_{\star}$. Thus, we can regard $y$ as a zero mode. To estimate the length over which this description is valid, let us consider (50) to leading order in $\beta_{I}$. After using the matching condition for $\Delta\theta$ one finds
$S_{\rm D3-tube}=\frac{m\beta_{1}}{2}\frac{1}{1+(y-y_{\star})}+\mathcal{O}(\beta_{1}^{2})\,.$ (52)
Thus, along a region of size $y-y_{\star}\lesssim\beta$, the denominator in (52) is approximated to one at leading order in $\beta_{1}$. Consequently, along a region of size $\sim\beta_{1}$, $y$ is indeed a flat direction. Hence, roughly speaking, the tube will contribute a prefactor
$Z_{\rm D3-tube}\sim\beta_{1}\,e^{-\frac{m\beta_{1}}{2}}\,.$ (53)
Since tube-nucleation is to be expected as $\beta_{1}\rightarrow 0$, for small $\beta_{1}$ we expect the contribution of the anchored geodesic to be
$e^{-2\widehat{S}_{\rm os}^{\rm(anchored)}}=Z_{\rm D3-tube}\,e^{-2S_{\rm os}^{\rm(anchored)}}\sim\frac{\beta_{1}\,e^{-\frac{m\beta_{1}}{2}}\,e^{-4m\beta_{2}}}{(X^{2}+Y^{2})^{m}}\,$, (54)
which then smooths the $\Theta(\beta_{1})$ contribution of the anchored geodesic. Note however that this result is proportional to $e^{-m\beta_{1}}$, so what really controls the anchored geodesic is the combination $m\beta_{1}$. This suggests that $\beta_{1}$ is at most of order $m^{-1}$ for this contribution to be non-negligible. Since $e^{-2\widehat{S}_{\rm os}^{\rm(anchored)}}\sim\beta_{1}$, we would have that this VEV is of order $m^{-1}$.

### v2

@@ 1 Introduction
As we will show, the problem of studying holographic correlators for giant gravitons boils down to studying charged geodesics hanging from the boundary in the 5d gauged SUGRA. This is essentially the WKB approximation to the two-point function [17] (see [18, 19] for applications to a context similar to ours). It turns out that, in the presence of the monodromy defect, there are two relevant such geodesics. One is the direct relative of the U-shaped geodesic capturing the two-point function in the absence of the defect. The other is a novel saddle corresponding to a geodesic anchored to the defect whose existence is tied to the presence of the defect. The full two-point function is the sum of both contributions, with the anchored contribution capturing the one-point function for the operator $OO^{\dagger}$ in the presence of the defect. It is important to emphasize that while the particular coefficients will depend on details of the defect, the structure and presence of the contribution itself is universal, only controlled by the existence of the geodesic and the asymptotics of the background.
@@ 2 Giant gravitons as charged geodesics in 5d gauged supergravity
Thus we see that effectively our D3 brane behaves like a charged particle, with charge equals to the mass, in the 5d geometry, precisely as expected for a giant graviton. [^3] We are assuming the appropriate boundary terms so as to impose fixed momentum. See [20, 21] for a more thorough discussion. The mass $m$ coincides with the dimension of the giant graviton and charge equaling the mass reflects the BPS condition saturated by the giant. We emphasize that (8) shows that $m\sim N$.
@@ 5 Comments on the anchored geodesics
It should be stressed that in the case of the anchored geodesic, this result does not rely on a perturbative expansion in $\beta_{1}$. From here it follows that, provided $\beta_{1}\neq 0$, the anchored geodesic looks like the cartoon in figure 1\@vpageref[]fig:anchored.
@@
On the other hand, this non-analytic behavior in a two-point function is strange from the standpoint of finite $N$ QFT. [^7] Note that the one-point function of square of the charged operator is not obviously controlled by a central charge and a priori depends on the continuous defect parameters. For instance, in a similar situation albeit for a free theory, the computation of [22] shows a smooth turning-off of the 1-point function for the square of the charged operator. Thus, it is natural to expect the existence of a mechanism to smooth this $\Theta$ behavior. One possible such mechanism is suggested by inspection of figure 1\@vpageref[]fig:anchored, since, as $\beta_{1}\rightarrow 0$ the configuration looks like a brane/antibrane prone to decay through nucleation of D3-brane tubes breaking the anchored geodesic. A cartoon of the configuration would be as in figure 2\@vpageref[]fig:instability below.
@@
$S_{\rm D3-tube}=m\frac{P^{\frac{1}{2}}}{H^{\frac{1}{3}}}\,\Delta\theta\,.$ (49)
@@
$S_{\rm D3-tube}=m\frac{(P^{\prime}_{\star})^{\frac{1}{2}}}{H_{\star}^{\frac{1}{3}}}\,\sqrt{y-y_{\star}}\,\Delta\theta\sim m\,\sqrt{y-y_{\star}}\,\Delta\theta\,.$ (50)
Since we want the tube to meet the two arms of the $U$, the $\Delta\theta$ should match that of the geodesic. In turn, using the equation of motion, we can estimate that for those $\Delta\theta\sim\beta\,\sqrt{y-y_{\star}}$. Hence
$S_{\rm D3-tube}\sim m\,\beta_{1}\,(y-y_{\star})\,.$ (51)
It should be stressed that this result does not rely on a perturbative expansion in $\beta_{1}$.
Thus we see that over a distance $y-y_{\star}\sim(m\beta_{1})^{-1}$, the contribution of the tube $e^{-S_{\rm D3-tube}}$ is much bigger than the contribution of the geodesic itself, which goes as $e^{-m}$. This shows that there is a “boundary layer” effect over distances $(m\beta_{1})^{-1}$ in which there could be important effects correcting the picture in terms of the geodesic followed by the D3 brane. [^8] It is straightforward to check that, for standard geodesics, the action for a similar tube action does not lead to a boundary layer. It is tempting to conjecture that properly taking into account this boundary layer leads to a smoothing of the non-analytic $\Theta(\beta_{1})$ behavior. In particular, such behavior might be an artifact of our geodesic approximation, which takes first the $m\rightarrow\infty$ limit –erasing the boundary layer– and only later $\beta_{1}\rightarrow 0$. Indeed, as stressed, for the heuristic argument for the existence of the boundary layer we are not relying on a perturbative expansion in $\beta_{1}$ –while we are implicitly assuming $m$ large by the mere use of a probe brane–, so one could generically estimate the boundary layer length as $m^{-1}$. This might generate a prefactor proportional to $m^{-1}$. On the other hand, the boundary layer thickness is really controlled by $m\beta_{1}$. We could then imagine, for instance, the double scaling limit where we send $\beta_{1}\rightarrow 0$ with fixed $m\beta_{1}$, so that the would-be $m^{-1}$ should be interpreted as $\beta_{1}/(m\beta_{1})\sim\beta_{1}$, thus suggesting that indeed the $\Theta$ is smoothed.